OpenStudy (anonymous):

Consider the system of inequalities: y>1/2x - 1 y <= -2x Which one of the following coordinates is a solution to the system? Select one: a. (0, 0) b. (-4, -4) c. (6, 6) d. (-7, 3)

OpenStudy (mathmate):

Let's see. I'll try by putting in each point into each inequality and see if they are satisfied. For (6,6), 6>(1/2)*6-1=2 6>2 so good. 6\(\le\)-2(6)=-12 but 6 is not less than nor equal to 12, so no good. (6,6) is NOT a choice. You can go through the same process for the other choices and see which one works.